# WikiDBGraph: A Data Management Benchmark Suite for Collaborative Learning over Database Silos thanks: *Equal Contribution

### v2

@@ VI-C Case 3: Multiple databases with Hybrid Overlap
VII Discussion
TABLE XII: Summary of key findings and implications for practical Collaborative Learning (CL)
Finding / Implication for Collaborative Learning Supporting Evidence (This Work) Reference(s)
CL overall outperforms Solo training on real-world databases. CL algorithms such as FedAvg, FedProx, and FedOV show improvements over Solo in both the overall analysis (Figs. 6, 5) and the case studies (Tables VIII, X). Aligns with [4, 46]
Advanced CL algorithms do not consistently outperform FedAvg on non-IID data. Relational schemas typically contain non-IID data, producing performance discrepancies across clients, as observed in the overall analysis (Figs. 6, 5) and the case study (Table VIII). Aligns with [4]
Databases in CL form highly interconnected networks with a long-tailed degree distribution; links should neither be ignored nor treated uniformly. WikiDBGraph exhibits a long-tailed node-degree distribution (Fig. 4(a)), with low median but high maximum degrees (Table IV). Similar to social/web networks [42, 43]; newly reported for CL
New CL paradigms are needed for relational databases to avoid computationally infeasible full-table joins. The “AllJoinSize” is infeasible on average ($9.5\times 10^{16}$ rows; Table V), making single-table approaches impractical. Novel finding
CL algorithms should handle hybrid cases with partial instance and feature overlap, beyond ideal alignment assumptions. Partial data and schema overlap are evidenced by the low OverlapRatio and JaccardColumn (Table VI) and highlighted in the hybrid case study (Section VI-C). Novel finding
Schema matching and entity resolution are critical, unresolved bottlenecks for real-world CL. Manual schema/feature alignment in case studies (Section VI) yields strong Combined performance, whereas in automated evaluation only a subset of CL tasks benefit and average improvements are small for both Combined and CL algorithms. Novel finding
This section summarizes key findings and implications for practical CL tasks derived from WikiDBGraph; a concise overview appears in Table XII. Some findings are consistent with established studies [4, 46], others parallel results reported in adjacent research areas [42, 43] but not previously for CL, and many are novel to our analysis. Collectively, these results highlight the unique complexities of real-world data collaboration and motivate further investigation.
Effectiveness of Collaborative Learning
Both the overall analysis (Figs. 6, 5) and the case studies (Tables VIII, X) show that CL algorithms generally outperform Solo training on real-world databases. However, advanced CL methods such as FedProx [47] do not consistently improve over simpler methods like FedAvg [46] on non-IID data. Both observations align with prior CL benchmarks [4], supporting WikiDBGraph as a valid benchmark for evaluating CL effectiveness.
Interconnection of Collaborative Learning
Databases in WikiDBGraph form highly interconnected networks with a long-tailed degree distribution. As evidenced by Fig. 4(a) and Table IV, some nodes have high degree while many have low degree. Thus, most existing CL approaches [44, 48] that ignore links or treat all nodes uniformly are suboptimal. This property mirrors patterns in other real-world networks (e.g., social and web graphs) [42, 43] and underscores the need for graph-aware CL frameworks.
Infeasibility of Full Table Joins.
The average AllJoinSize—the estimated row count from joining all tables within a database—is $9\times 10^{16}$ (Table V), rendering single-table approaches computationally infeasible. Consequently, CL methods that require full table joins are impractical for large real-world databases. Although such databases may be less common, they are often highly valuable due to their scale, calling for CL frameworks that operate directly on relational structures rather than a single prejoined table.
Hybrid Alignment in Collaborative Learning
Databases in WikiDBGraph are neither purely horizontally nor purely vertically aligned; instead, they exhibit hybrid alignment. This is supported by low OverlapRatio and JaccardColumn values in Table VI and is further elaborated in the hybrid case study (Section VI-C). As a result, CL approaches that assume ideal horizontal or vertical alignment cannot handle many real-world cases, motivating methods that explicitly accommodate hybrid alignment.
VIII Future Directions
This section presents three main opportunities for data management research on WikiDBGraph.
Schema Matching: Analyses on WikiDBGraph indicate strong potential for schema matching [50], as identifying correct feature correspondences is crucial for learning accuracy. The dataset also provides ground truth for schema matching not only between two databases but also across multiple databases with complex relationships. WikiDBGraph surpasses prior datasets such as GitSchema [32] in scale and offers rich data beyond schemas alone. Its data distributions can further aid column alignment, enabling data-aware schema matching approaches.
Record Linkage and Data Mining: The prevalence of unaligned data suggests that record linkage [51] (a.k.a. entity alignment [52]) is inherently noisy [53]. Improving alignment in isolation, without feedback from downstream data mining, hinders task-aware alignment. A critical direction is to mine data under inaccurate or incomplete row alignment, which typically calls for a coupled data-mining framework [54].
Graph Mining: To incorporate databases in CL that cannot be directly joined, graph mining [55] techniques must be extended to support CL over database graphs—tables connected by foreign keys. This join challenge calls for new graph-based CL methods beyond current table-centric approaches.
IX Conclusion
@@
This paper presented WikiDBGraph, a large-scale graph of 100,000 interconnected databases designed to address the critical lack of inter-database relationships in existing corpora. We demonstrated that WikiDBGraph successfully identifies correlated databases that benefit from collaborative learning, while also highlighting the performance gap compared to centralized models and surfacing novel challenges. Crucially, our analysis pinpoints data preprocessing, rather than the learning algorithms themselves, as the primary bottleneck in practical deployments. WikiDBGraph thus serves as a crucial benchmark suite to measure this gap and guide future research into more effective and scalable collaborative learning paradigms.

### v3

@@ VII-D Case 4: Multiple databases with Hybrid Overlap
VIII Conclusion
@@
This paper presented WikiDBGraph, a large-scale graph of 100,000 interconnected databases designed to address the critical lack of inter-database relationships in existing corpora. We demonstrated that WikiDBGraph successfully identifies correlated databases that benefit from collaborative learning, while also highlighting the performance gap compared to centralized models and surfacing novel challenges. WikiDBGraph thus serves as a crucial benchmark to measure this gap and guide future research into more effective and scalable collaborative learning paradigms.
@@
Acknowledgement
This research/project is supported by the National Research Foundation, Singapore and Infocomm Media Development Authority under its Trust Tech Funding Initiative. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore and Infocomm Media Development Authority.